Step by step
Original- ax^2 + bx + c =0
1. subtract c from each side
2. Divide each side by a
3. Add he square of half the coefficient of x to each side
4.write the left side as a perfect square
5.use a common denominator to express the right side as a single fraction
6. find the square root of eac side
7. solve for x by subtracting the same term form each side
8. use a common denominator to express the right side as a single fraction
once all the steps are done you should end up with the quadratice formula but i cant get past step three and i need each step. PLEASE HELp

Answer by venugopalramana(3286)

Step by step
Original- ax^2 + bx + c =0
1. subtract c from each side...OK...WE GET AX^2+BX=-C
2. Divide each side by a......OK...WE GET..X^2+BX/A=-C/A
3. Add he square of half the coefficient of x to each side..OK..WE GET
X^2+BX/A+(B/2A)^2=(B/2A)^2-(C/A)
4.write the left side as a perfect square...OK...WE GET ..
{X+(B/2A)}^2=B^2/4A^2 - C/A
5.use a common denominator to express the right side as a single fraction...OK...
{X+(B/2A)}^2=(B^2-4AC)/4A^2
6. find the square root of eac side....OK......WE GET
X+(B/2A)=+ OR - {SQRT(B^2-4AC)}/2A
7. solve for x by subtracting the same term form each side...OK....
X=-(B/2A)+OR-{SQRT(B^2-4AC)}/2A
8. use a common denominator to express the right side as a single fraction
X=[-B+OR-{SQRT(B^2-4AC)}]/(2A)
THAT IS IT GOT IT?..OK...
once all the steps are done you should end up with the quadratice formula but i cant get past step three and i need each step. PLEASE HELp